# Parametric Mean-Field Empirical Bayes in High-dimensional Linear Regression

### v1

@@ A.3 Proofs from Section 4
The next technical result provides a way to quantify the distance between the true posterior and the EB posterior. As it involves the use of several technical lemmas from [34] to compute moment bounds for RFIMs, we state them in the following lemma. Each part follows from Theorem 2.3, Lemma 3.2(b), and Lemma 3.1 in [34], respectively. We also refer the reader to [7, Lemma 2,10] and [17, Lemma A.1] for related results.
@@
    For a constant vector $\gamma\in\mathbb{R}^{p}$, we have $\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\gamma_{i}(B_{i}-u_{i,\theta})\Big]^{2}=\mathnormal{O}_{\mathnormal{P},X}\big(\|\gamma\|^{2}\big).$
@@
As the original result in [34] allows any choice of prior/base-measure $\mu$, it also holds under the estimated prior $\mu=\mu_{\hat{\theta}_{\mathnormal{p}}}$. Here, we have slightly modified the statement in [34] (i.e. replaced an additional error term $R_{3p}$ there in terms of $R_{2}$) by replacing the bound in the third display in [pg 17, 34] to the following:
@@
    $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\Big[\psi_{0,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})-\psi_{0,\theta_{0}}^{\prime}(w_{i})\Big]=(\hat{\theta}_{\mathnormal{p}}-\theta_{0})^{\top}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\frac{\partial g}{\partial\theta}(\theta_{0},w_{i})+o_{P}(1)\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,\gamma^{2}J(\theta_{0})^{\top}S(\theta_{0})J(\theta_{0})\Big),$
@@ B.1 Proof of Lemmas from the main paper
Here the first equality used the first Bartlett’s identity $\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{\theta})=0$, and the second equality follows from the second Bartlett’s identity $I(\theta)=-\operatorname{\mathbb{E}}\big[\nabla^{2}\ell(\theta;B_{\theta})\big]=\operatorname{\textnormal{Var}}\nabla\ell(\theta;B_{\theta})$. ∎
@@
    As this is a continuous function (of $d,\theta$) on a compact domain, it is bounded. For example, regarding the supremum over $d$, direct computations show that
@@
    R5 is immediate by applying Lemma B.1, as each mixture component of $B_{\theta_{0}}$ is sub-Gaussian by parts (a) and (b). The first part of R6 is also bounded by the tower property. To check the second part of R6, let $p_{k}(b)$ denote the conditional density (if we consider a degenerate mixture component, consider the base measure $\nu=\delta_{0}+\lambda$ where $\lambda$ denotes the Lebesgue measure) of the $k$th mixture component, and observe that $B_{t,d,\theta}$ has density:
    $\operatorname{\mathbb{P}}(B_{t,d,\theta}=b)\propto\sum_{k=1}^{K}\pi_{k}p_{k}(b)e^{tb-\frac{db^{2}}{2}}.$
    This is a finite mixture of the component-wise tilts $B^{(k)}_{t,d}$, which have density $p_{k}(b)e^{tb-\frac{db^{2}}{2}}$. By parts (a) and (b), each $B^{(k)}_{t,d}$ are sub-Gaussian with $\|B^{(k)}_{t,d}\|_{\psi_{2}}\leq K$ for some constant $K>0$. Also, using the properties of exponential tilting (in particular, we appropriately modify (54) in Lemma A.5), we can show that $\operatorname{\mathbb{E}}B^{(k)}_{t,d}=O(1+K|t|)$. Now, the conclusion follows by applying Lemma B.1.
@@ B.2 Proof of upper bound lemmas
The conclusion holds if the final inequality is strict for any $\theta\neq\theta_{0}$. This is because the parametric family $\{\operatorname{\mathbb{P}}_{W_{\theta}}\}_{\theta\in\Theta}$ (see (43)) is identifiable (i.e. $\operatorname{\mathbb{P}}_{W_{\theta}}=\operatorname{\mathbb{P}}_{W_{\theta^{\prime}}}$ implies $\theta=\theta^{\prime}$), which follows from the injectivity of the Laplace transform for finite measures, alongside the identifiability of $\mu_{\theta}$ from R0 in 3.2. ∎
@@
    Step 1. Recall that $w\mid X,\beta^{\star}\sim N(\Sigma\beta^{\star},\Sigma)$ where $\Sigma=\sigma^{-2}X^{\top}X$. By the Gaussian Poincare inequality [e.g. Theorem 3.25, 57], the growth condition, and Lemma A.7, we have
    $\operatorname{\textnormal{Var}}\Big(\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i})\mid X,\beta^{\star}\Big)\leq$ $\frac{1}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}[g^{\prime}(w_{i})^{2}\mid X,\beta^{\star}]\leq\frac{1}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[1+\operatorname{\mathbb{E}}\big[|w_{i}|^{2r}\mid X,\beta^{\star}\big]\Big]=O_{P}\Big(\frac{1}{p}\Big).$
@@
$\operatorname{\mathbb{E}}_{\beta_{i}^{\star}\sim B_{\theta_{0}}}\Phi(d_{0}\beta_{i}^{\star})=\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{\theta_{0}})=0,\quad\operatorname{\textnormal{Var}}_{\beta_{i}^{\star}\sim B_{\theta_{0}}}\big[\Phi(d_{0}\beta_{i}^{\star})\big]=\operatorname{\textnormal{Var}}\Big[\operatorname{\mathbb{E}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]=T_{1}.$
@@
$\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\textnormal{Var}}(h(w_{i}))=\underbrace{\operatorname{\mathbb{E}}\Big[\operatorname{\textnormal{Var}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]}_{=T_{2}}+o_{P,X}(1),\quad\frac{1}{p}\sum_{1\mathnormal{\leq i\neq j\leq p}}\operatorname{\textnormal{Cov}}(h(w_{i}),h(w_{j}))=o_{P,X}(1).$
@@ B.3 Proof of lower bound lemmas
    $H_{1}=\frac{1}{p\sqrt{p}}\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\gamma_{j}(B_{j}-u_{j}),$
    where $\gamma_{j}:=\sum_{i\neq j}A_{p}(i,j)h_{i}^{\prime}(s_{i}+c_{i})$. Then, by part (a) of Lemma A.10, we have
    $p^{2}\operatorname{\mathbb{E}}H_{1}^{2}\lesssim\frac{1}{p}\big(\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\gamma_{j}^{2}\big)\Big(1+\frac{p^{3}}{n^{2}}\Big)=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p}{n}\Big).$
    The final inequality follows from using the operator norm bound for $A_{p}$ in Lemma A.11 to write $\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\gamma_{j}^{2}\lesssim p\|A_{p}\|^{2}=\mathnormal{O}_{\mathnormal{P},X}(p^{2}/n)$.
@@
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[h_{i}(s_{i}+w_{i})-h_{i}(w_{i})\Big]$ $=\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[s_{i}h_{i}^{\prime}(\xi_{i}+w_{i})\Big]\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)s_{i}h_{i}^{\prime}(\xi_{i}+w_{i})u_{j}=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p\sqrt{p}}{\sqrt{n}}\Big).$ (59)
@@ B.4 Proof of lemmas for Section 4
$~~\big|\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)q_{i}(\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w_{i})-\hat{\upsilon}_{p})\psi_{j,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{j})\big|$
@@ Appendix C Proof of technical concentration lemmas
Assuming $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}(\beta_{1}^{\star})^{4}<\infty$ and $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}g(\beta_{1}^{\star})^{2}<\infty$, it holds that
@@
For notational convenience, define a $n\times p$ matrix $Z$ with entries $Z_{k,i}:=\sqrt{n}X_{k,i}$, so that each $Z_{k,i}$ are independent sub-Gaussians with mean 0 and variance 1. We will write $\operatorname{\mathbb{E}}_{\operatorname{Z}}$ and $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}$ for expectations with respect to the matrix $Z$ and $\operatorname{\boldsymbol{\beta}^{\star}}$. Observe that
@@
$\frac{n^{2}}{p^{4}}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}\operatorname{\textnormal{Var}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\boldsymbol{\beta}^{\star}}\right)\to 0\quad\mbox{and}\quad\frac{n^{2}}{p^{4}}\operatorname{\textnormal{Var}}\operatorname{\mathbb{E}}_{\operatorname{Z}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\boldsymbol{\beta}^{\star}}\right)\to 0.$ (65)
@@
$\;\;\;\;\frac{n^{2}}{p^{4}}\operatorname{\textnormal{Var}}\operatorname{\mathbb{E}}_{\operatorname{Z}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\boldsymbol{\beta}^{\star}}\right)$
@@
$\;\;\;\;\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}\operatorname{\textnormal{Var}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\boldsymbol{\beta}^{\star}}\right)$

### v2

@@ A.3 Proofs from Section 4
The next technical result provides a way to quantify the distance between the true posterior and the EB posterior. As it involves the use of several technical lemmas from [36] to compute moment bounds for RFIMs, we state them in the following lemma. Each part follows from Theorem 2.3, Lemma 3.2(b), and Lemma 3.1 in [36], respectively. We also refer the reader to [7, Lemma 2,10] and [18, Lemma A.1] for related results.
@@
    For a constant vector $\mathfrak{\zeta}\in\mathbb{R}^{p}$, we have $\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\zeta_{i}(B_{i}-u_{i,\theta})\Big]^{2}=\mathnormal{O}_{\mathnormal{P},X}\big(\|\zeta\|^{2}\big).$
@@
As the original result in [36] allows any choice of prior/base-measure $\mu$, it also holds under the estimated prior $\mu=\mu_{\hat{\theta}_{\mathnormal{p}}}$. Here, we have slightly modified the statement in [36] (i.e. replaced an additional error term $R_{3p}$ there in terms of $R_{2}$) by replacing the bound in the third display in [pg 17, 36] to the following:
@@
    $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\Big[\psi_{0,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{i})-\psi_{0,\theta_{0}}^{\prime}(w_{i})\Big]$ $=(\hat{\theta}_{\mathnormal{p}}-\theta_{0})^{\top}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\frac{\partial g}{\partial\theta}(\theta_{0},w_{i})+o_{P}(1)$
    $\operatorname*{\xrightarrow{\mathnormal{d}}}N\Big(0,\gamma^{2}J(\theta_{0})^{\top}S(\theta_{0})J(\theta_{0})\Big),$
@@ B.1 Proof of Lemmas from the main paper
Here the first equality used the first Bartlett’s identity $\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{\theta})=0$, and the second equality follows from the second Bartlett’s identity $I(\theta)=-\operatorname{\mathbb{E}}\big[\nabla^{2}\ell(\theta;B_{\theta})\big]=\operatorname{\textnormal{Var}}\nabla\ell(\theta;B_{\theta})$ (see R3). ∎
@@
    As this is a continuous function (of $d,\theta$) on a compact domain (by R1), it is bounded. For example, regarding the supremum over $d$, direct computations show that
@@
    R5 is immediate by applying Lemma B.1, as each mixture component of $B_{\theta_{0}}$ is sub-Gaussian by parts (a) and (b). The first part of R6 is also bounded by the tower property. To check the second part of R6, let $p_{m}(b)$ denote the conditional density (if we consider a degenerate mixture component, consider the base measure $\nu=\delta_{0}+\lambda$ where $\lambda$ denotes the Lebesgue measure) of the $m$th mixture component, and observe that $B_{t,d,\theta}$ has density:
    $\operatorname{\mathbb{P}}(B_{t,d,\theta}=b)\propto\sum_{m=1}^{M}\pi_{m}p_{m}(b)e^{tb-\frac{db^{2}}{2}}.$
    This is a finite mixture of the component-wise tilts $B^{(m)}_{t,d}$, which have density $p_{m}(b)e^{tb-\frac{db^{2}}{2}}$. By parts (a) and (b), each $B^{(m)}_{t,d}$ are sub-Gaussian with $\|B^{(m)}_{t,d}\|_{\psi_{2}}\leq K$ for some constant $K>0$. Also, using the properties of exponential tilting (in particular, we appropriately modify (55) in Lemma A.5), we can show that $\operatorname{\mathbb{E}}B^{(m)}_{t,d}=O(1+K|t|)$. Now, the conclusion follows by applying Lemma B.1.
@@ B.2 Proof of upper bound lemmas
The conclusion holds if the final inequality is strict for any $\theta\neq\theta_{0}$. This is because the parametric family $\{\operatorname{\mathbb{P}}_{W_{\theta}}\}_{\theta\in\Theta}$ (see (44)) is identifiable (i.e., $\operatorname{\mathbb{P}}_{W_{\theta}}=\operatorname{\mathbb{P}}_{W_{\theta^{\prime}}}$ implies $\theta=\theta^{\prime}$), which follows from the injectivity of the Laplace transform for finite measures, alongside the identifiability of $\mu_{\theta}$ from R0 in 3.2. ∎
@@
    Step 1. Recall that $w\mid X,\beta^{\star}\sim N(\Sigma\beta^{\star},\Sigma)$ where $\Sigma=\sigma^{-2}X^{\top}X$. By the Gaussian Poincare inequality [e.g. Theorem 3.25, 60], the growth condition, and Lemma A.7, we have
    $\operatorname{\textnormal{Var}}\Big(\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i})\mid X,\beta^{\star}\Big)$ $\leq\frac{1}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}[g^{\prime}(w_{i})^{2}\mid X,\beta^{\star}]$
    $\leq\frac{1}{p^{2}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[1+\operatorname{\mathbb{E}}\big[|w_{i}|^{2r}\mid X,\beta^{\star}\big]\Big]=O_{P}\Big(\frac{1}{p}\Big).$
@@
$\operatorname{\mathbb{E}}_{\beta_{i}^{\star}\sim B_{\theta_{0}}}\Phi(d_{0}\beta_{i}^{\star})=\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{\theta_{0}})=0,~\operatorname{\textnormal{Var}}_{\beta_{i}^{\star}\sim B_{\theta_{0}}}\big[\Phi(d_{0}\beta_{i}^{\star})\big]=\operatorname{\textnormal{Var}}\Big[\operatorname{\mathbb{E}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]=T_{1}.$
@@
$\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\textnormal{Var}}(h(w_{i}))=\underbrace{\operatorname{\mathbb{E}}\Big[\operatorname{\textnormal{Var}}\big[h(W_{\theta_{0}})\mid B_{\theta_{0}}\big]\Big]}_{=T_{2}}+o_{P,X}(1),~\frac{1}{p}\sum_{1\mathnormal{\leq i\neq j\leq p}}\operatorname{\textnormal{Cov}}(h(w_{i}),h(w_{j}))=o_{P,X}(1).$
@@ B.3 Proof of lower bound lemmas
    $H_{1}=\frac{1}{p\sqrt{p}}\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\zeta_{j}(B_{j}-u_{j}),$
    where $\zeta_{j}:=\sum_{i\neq j}A_{p}(i,j)h_{i}^{\prime}(s_{i}+c_{i})$. Then, by part (a) of Lemma A.10, we have
    $p^{2}\operatorname{\mathbb{E}}H_{1}^{2}\lesssim\frac{1}{p}\big(\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\zeta_{j}^{2}\big)\Big(1+\frac{p^{3}}{n^{2}}\Big)=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p}{n}\Big).$
    The final inequality follows from using the operator norm bound for $A_{p}$ in Lemma A.11 to write $\sum_{\mathnormal{j}=1}^{\mathnormal{p}}\zeta_{j}^{2}\lesssim p\|A_{p}\|^{2}=\mathnormal{O}_{\mathnormal{P},X}(p^{2}/n)$.
@@
$\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[h_{i}(s_{i}+w_{i})-h_{i}(w_{i})\Big]$
$=$ $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Big[s_{i}h_{i}^{\prime}(\xi_{i}+w_{i})\Big]\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)s_{i}h_{i}^{\prime}(\xi_{i}+w_{i})u_{j}=\mathnormal{O}_{\mathnormal{P},X}\Big(\frac{p\sqrt{p}}{\sqrt{n}}\Big).$
@@ B.4 Proof of lemmas for Section 4
$\big|\sum_{1\mathnormal{\leq i\neq j\leq p}}A_{p}(i,j)q_{i}(\psi_{i,\hat{\theta}_{\mathnormal{p}}}^{\prime\prime}(w_{i})-\hat{\upsilon}_{p})\psi_{j,\hat{\theta}_{\mathnormal{p}}}^{\prime}(w_{j})\big|$
@@ Appendix C Proof of technical concentration lemmas
Assume $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}(\beta_{1}^{\star})^{4}<\infty$ and let $g$ be a function such that $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}g(\beta_{1}^{\star})^{2}<\infty$. Then, it holds that
@@
Note that the moment requirements in Lemma C.1 are immediate by R5.
@@
For notational convenience, define a $n\times p$ matrix $Z$ with entries $Z_{k,i}:=\sqrt{n}X_{k,i}$, so that each $Z_{k,i}$ are independent sub-Gaussians with mean 0 and variance 1. We will write $\operatorname{\mathbb{E}}_{\operatorname{Z}}$ and $\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}$ for expectations with respect to the matrix $Z$ and $\operatorname{\beta^{\star}}$. Observe that
@@
$\frac{n^{2}}{p^{4}}\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}\operatorname{\textnormal{Var}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\beta^{\star}}\right)\to 0\quad\mbox{and}\quad\frac{n^{2}}{p^{4}}\operatorname{\textnormal{Var}}\operatorname{\mathbb{E}}_{\operatorname{Z}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\beta^{\star}}\right)\to 0.$ (65)
@@
$\;\;\;\;\frac{n^{2}}{p^{4}}\operatorname{\textnormal{Var}}\operatorname{\mathbb{E}}_{\operatorname{Z}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\beta^{\star}}\right)$
@@
$\;\;\;\;\operatorname{\operatorname{\mathbb{E}}_{\beta^{\star}}}\operatorname{\textnormal{Var}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\mathcal{F}_{i}^{2}g(\beta_{i}^{\star})\mid\operatorname{\beta^{\star}}\right)$